# NuiScene: Exploring Efficient Generation of Unbounded Outdoor Scenes

### v1

@@ 4.1 Dataset
For single-scene experiments, we use the processed scene from Figure 5 and follow the chunk sampling process outlined in Sec. 3.1 for VAE training. Instead of sampling individual chunks, we sample larger quad-chunks of size $(100,h_{vox},100)$, which consists of a 2x2 grid of smaller chunks $(50,h_{vox},50)$. Sampling is done along the $x$ and $z$ axes using occ[i-50:i+50,:,k-50:k+50]. We sample 100k quad-chunks in total, with 95k used for training and 5k for validation. While these chunks may overlap, they are not identical, as they come from different $x$ and $z$ coordinates. We define the batch size as the number of quad-chunks used for training, with the quad-chunks divided into 4 smaller chunks during VAE training. For the diffusion model, the quad-chunks’ point clouds are encoded using the VAE for diffusion training with the 4 different patterns.
For multi-scene experiments, we add three additional scenes in additional to the single scene for training. In total we sample 300k quad-chunks from these scenes. Please see the supplement (Figure 10) for more visualizations.
@@ 4.2 Evaluation Metrics
Reconstruction For reconstruction metrics, we report Chamfer Distance (CD), F-Score, and IoU following Zhang et al. [49] by comparing ground truth and reconstructed chunks on the validation set. CD and F-Score are computed from 50k sampled points per chunk using a threshold of $2/50$, the voxel length of our sampled chunks. For IoU, we sample 20k occupancy values, with 5k from occupied regions and 5k from unoccupied regions, along with 10k near-surface points along the chunk’s surface.
@@
Generation We evaluate the generation quality of our diffusion models using Fréchet PointNet++ [34] Distance (FPD) and Kernel PointNet++ Distance (KPD), following Zhang et al. [49]. We use the Pointnet++ model from Point-E [33] following Yan et al. [47]. Specifically, we sample 10k quad-chunks from the diffusion model, compute their surface point clouds (2048 points), and compare them to 10k sampled quad-chunks from the original scene.
@@
Table 1: Comparison of VAE training resources for vector set vs triplane backbones. Training for all experiments was run on 2 L40S GPUs, total batch size and memory across 2 gpus are reported. The # Latents is the size of the latent for the VAE backbone and Output Res indicates the triplane size after deconvolution. Increasing the output triplane resolution to $3\times 128^{2}$ causes the model to exceed memory limits (OOM).
@@
triplane 40 $3\times 4^{2}$ $3\times 32^{2}$ 6 35.9 35.7
@@ 40 $3\times 4^{2}$ $3\times 128^{2}$ 6 - OOM
4.3 Model and Baselines
We use triplane representations as a baseline for this task, given their widespread use in prior work [26, 44]. We do not directly adopt BlockFusion’s triplane architecture, as it requires fitting triplanes for all dataset chunks, a process that took thousands of GPU hours on 3D-Front. Instead, we implement a VAE backbone with architecture similar to LRM[15] discussed in Sec. 3.2.
We exclude dense grid methods [30, 32] due to their high memory demands, which scale cubically with resolution, compared to the quadratic growth of triplanes. In our task, the height can be up to about 8.5 times greater than the $x$ and $z$ dimensions for scene chunks in Figure 5, further exacerbating memory usage. Additionally, LT3SD [32] requires multi-scale training for both the autoencoder and diffusion models, increasing training time.
VAE For triplane baselines, we explore configurations with $V=3\times 4\times 4$, varying the number of deconvolution layers, which affects output resolution, and the normalization scale $S$ (see Tab. 2). For the vector set model, compressing to $V=16$ preserves geometric fidelity while enabling efficient diffusion training. Both representations use a channel size of $c=64$, and a sampled point cloud size of $N_{p}=4096$. We set the output triplane channel $c_{tri}=40$. The decoder has $L=24$ self-attention layers, and all models are trained for $160$ epochs. Each training iteration samples $4096$ query points per chunk for occupancy supervision.
Diffusion For all models, we use a 25-layer UNet-like transformer [28] for diffusion. The triplane diffusion model operates on embeddings from the VAE with the output triplane resolution of $h_{tri}=w_{tri}=64$ and $S=6$. We train all models for $320$ epochs and a batch size of $192$. The transformer processes tokens of length $4\times V$ as the diffusion model generates a 2x2 grid of chunks at once.
4.4 Single Scene
In single-scene experiments, we demonstrate the effectiveness of representing scene chunks as vector sets compared to spatially structured latents like triplanes. VAE reconstruction results (Sec. 4.4.1) and diffusion model comparisons (Sec. 4.4.2) show that the vector set representation achieves better compression, leading to improved training efficiency and superior performance over triplanes.
4.4.1 VAE Reconstruction
Tab. 3 presents the quantitative reconstruction results, showing that the vector set model outperforms all triplane settings. We also report whether occupancy queries are pruned using the predicted height $\hat{h}$ or the ground truth height $\tilde{h}$ when computing meshes for CD and F-Score. The similar values in both cases indicate the accuracy of our height prediction.
The triplane models are constrained by the VAE’s output resolution, with performance improving as resolution increases from $3\times 32^{2}$ to $3\times 64^{2}$. However, for tall scenes, coordinate clamping during triplane sampling prevents reconstruction of tall buildings exceeding the scaling factor of $S=6$. Increasing $S$ to $8.5$ (matching the tallest chunk) severely degrades performance due to precision loss from excessive compression along the y-axis. While further increasing resolution to $128^{2}$ via deconvolution is possible, it results in out-of-memory (OOM) on two L40S GPUs ( Tab. 2). In contrast, the vector set model provides better compression ($V=16$ vs $V=3\times 4\times 4$) while requiring similar training time and memory as the lowest-resolution triplane model, demonstrating its efficiency in representing scene chunks and superior performance.
@@ 4.4.2 Diffusion
Figure 7: We compare generation results using RePaint [31] for outpainting for resampling steps $r=1$ and $r=5$.
Figure 8: We present results from our NuiScene model trained on multiple scenes. In Figure (a), the top-right shows the generated geometry, while the top-left displays the textured model using SceneTex [6]. We render these models and feed them into Trellis [45] for reconstruction (zoom in to see details). It can be seen that our method showcases better geometric detail. Figure (b) presents a large scene generated by our model, showcasing its ability to blend elements from different scenes in the dataset.
Quality. The FPD and KPD scores (Tab. 4) show that the vector set diffusion model outperforms the baseline triplane diffusion model as well. Qualitative results for a 21x21 chunk scene (Figure 6) demonstrate the vector set model’s ability to generate finer details, especially for buildings, whereas the triplane model struggles with details and introduces noisy artifacts which contributes to the higher FPD and KPD scores, likely due to compression of query points along the y axis. Additionally, as shown in Tab. 2, the vector set model trains about 2.5 times faster and uses half the VRAM, benefiting from three times fewer tokens with a smaller $V$.
Table 5: We benchmark generation time on 1 RTX 2080Ti for generating $21\times 21$ chunks using RePaint with 5 resampling steps and our explicit outpainting method with the vecset diffusion model, and also report the triplane diffusion model’s generation time.
@@
vecset RePaint $21\times 21$ 1030.54 60.74
triplane explicit $21\times 21$ 219.42 30.21
vecset explicit $21\times 21$ 217.79 76.59
Outpainting. We compare against the inpainting method RePaint [31], widely used in prior works [44, 32, 26] for unbounded outpainting. To ensure a fair comparison, we generate scenes in the same order and with the same overlaps as our outpainting model. However, in RePaint, the mask and conditional embeddings are set to zero, with existing chunk conditions incorporated through its iterative re-sampling process. RePaint results for resampling steps $r=1$ and $r=5$ are shown in Figure 7. Without resampling ($r=1$), inter-chunk connections lack coherence, and the process sometimes collapses into broken chunks, likely due to the model repeatedly conditioning on noisy diffusion-generated chunks. With $r=5$, coherence improves, but the RePaint baseline still produces less diverse scenes and occasional collapses, even with more resampling steps (Figure 14, supplement). Additionally, iterative re-sampling is significantly slower than our outpainting method (Tab. 5), as inference time scales with the number of resampling steps, whereas our method does not require resampling.
Our outpainting method is not perfect, and some chunks may still be poorly connected or discontinuous. In Figure 6, the right orange box shows coherently connected chunks for our vector set diffusion model, while the left-side orange box highlights issues such as misaligned bridges or non-continuous chunks with visible gaps and seams. We believe that this is largely due to the limited dataset size, where certain objects may not appear often enough, leading to insufficient training examples for seamless generation.
@@ 4.5 Multiple Scenes
In Sec. 4.4, we demonstrated the advantages of vector sets over triplanes for efficient representation, allowing flexible handling of scene chunks with varying heights. As well as the effectiveness of our outpainting model for fast large scale outdoor scene generation. However, training on a single scene leads to overfitting, with sub-scenes resembling the training data, despite the model’s ability to generate some novel compositions ( Figure 13, supplement). To assess its generalization ability and validate our curation process for NuiScene43, we extend training to three additional scenes using the vector set backbone.
Figure 8 (b) shows a scene generated by our model, with blue boxes highlighting how it integrates elements (rural houses, aqueduct, skyscraper) from two distinct scenes. This demonstrates its generalization ability, as such combinations are absent from the training data. Our model’s efficient generation process, combined with a curation process that unifies diverse scenes into a cohesive training dataset, enables the seamless stitching of chunks from disjoint environments.
We further compare our approach with Trellis [45], a state-of-the-art image-to-3D object generation method, using its official HuggingFace demo. The top right of Figure 8 (a) shows the geometry produced by our model, with the top left displaying a textured version using SceneTex. We render both textured and geometry-only meshes and use them as inputs for Trellis, with results shown at the bottom. As seen, Trellis produces coarser geometric details, as its single latent representation for entire scenes limits fine details in large environments. In contrast, our model represents each chunk separately, enabling higher-detail and unbounded scene generation. This highlights a key limitation of current object generation methods, which struggle to scale effectively for unbounded scene generation.
@@ 5 Conclusion
Limitations One limitation of our model is the limited number of training scenes, as we currently pre-sample chunks offline, leading to high storage costs when sampling a large amount of chunks from scenes. More efficient scene representations, such as octrees, could enable online sampling and support training on the entire NuiScene43 dataset. Another limitation is the lack of control in our generative model, as we currently lack text descriptions or attributes for individual chunks, making controllable conditional generation difficult. However, we may be able to leverage foundation models [21, 1] to allow for semantic map or text conditioning. A related challenge, also present in previous works [44, 32, 26, 30], is the lack of global context and larger control. The model only conditions on a small portion of chunks for generation, making it difficult to make large-scale decisions, such as city planning with roads.
We introduce NuiScene, a method for expansive outdoor scene generation across heterogeneous scenes. Our approach addresses key challenges posed in this task, such as large height variations between scene chunks, by representing them as vector sets. This allows for efficient training and flexible querying, overcoming the limitations of prior works that rely on spatially structured latents. Another crucial aspect of large-scale outdoor generation is speed. Instead of slow inpainting methods like RePaint, we train an outpainting model that enables fast inference without resampling steps while ensuring better coherence between chunks. To support this task, we curate a high-quality dataset from moderate to large scenes, enabling the exploration of (1) varied scene heights and (2) heterogeneous scene blending. Trained on multiple scenes, our model showcases cross-scene generation by seamlessly merging distinct elements that never co-occurred in the training data, creating a cohesive scene and validating the effectiveness of our curation pipeline. We believe further automation of this process will expand available datasets, unlocking open-domain large-scale scene generation.
@@
Figure 9: Sample and depth mean deviation maps are calculated for sampling chunks from scenes. The sample map is a binary map defining valid x and z coordinates where quad chunks are fully occupied. And the depth mean variation map is used to filter out overly flat regions.
@@
Figure 10: Three additional scenes used to train our multi-scene model. The top two sub-scenes were split from the same scene in Objaverse for occupancy calculation. The shown scenes have been processed with fixed ground geometries, and their meshes were extracted using the marching cubes algorithm on the occupancy grid.
Appendix S1 NuiScene43 Dataset
@@
We obtain training scenes from Objaverse by applying several rounds of filtering, ultimately selecting 43 high quality moderately to large sized scenes (S1.1). To ensure a unified scale between scenes, we annotate scenes with scales relative to each other, aligning them under a uniform scale (S1.2). Finally, we preprocess the raw Objaverse mesh files for training by sampling point clouds, adjusting ground geometry, and converting them to occupancies (S1.3). Furthermore, we describe the sampling maps used to sample quad chunks from scenes (S1.4) as well as some dataset statistics (S1.5).
S1.1 Filtering
@@
We begin by filtering Objaverse [9] using multi-view embeddings from DuoduoCLIP [24]. We use cosine similarity to query embeddings and retrieve objects, then refine selection through manual labeling and neural network filtering trained on DuoduoCLIP embeddings, reducing 37k initial scenes to 2k. Since manually processing all scenes for scale labeling and ground alignment is impractical, we select 43 larger scenes for initial experiments, which we refer to as the NuiScene43 dataset.
S1.2 Scale Labeling

### v2

@@ 4.1 Dataset
For single-scene experiments, we use the processed scene from Figure˜5 for training. Instead of sampling individual chunks, we sample larger quad-chunks of size $(100,h_{vox},100)$, which consists of a 2x2 grid of smaller chunks $(50,h_{vox},50)$. Sampling is done along the $x$ and $z$ axes using occ[i-50:i+50,:,k-50:k+50]. We sample 100k quad-chunks in total, with 95k used for training and 5k for validation. While these quad-chunks may overlap, they are not identical, as they come from different $x$ and $z$ coordinates. We define the batch size as the number of quad-chunks used for training, with the quad-chunks divided into 4 smaller chunks during VAE training (380k train/20k val). For the diffusion model, the quad-chunks’ point clouds are encoded using the VAE for diffusion training.
For multi-scene experiments, we add three additional scenes in additional to the single scene for training. In total we sample 300k quad-chunks from these scenes.
@@ 4.2 Evaluation Metrics
Reconstruction. For reconstruction metrics, we report Chamfer Distance (CD), F-Score, and IoU following Zhang et al. [52] by comparing ground truth and reconstructed chunks on the validation set. CD and F-Score are computed from 50k sampled points per chunk using a threshold of $2/50$ (the voxel length of sampled chunks). For IoU, we sample 20k occupancy values, with 10k split between occupied and unoccupied regions, along with 10k near-surface points.
@@
Generation. We evaluate the generation quality of our diffusion models using Fréchet PointNet++ [36] Distance (FPD) and Kernel PointNet++ Distance (KPD), following Zhang et al. [52]. We use the Pointnet++ model from Point-E [34] following Yan et al. [50]. Specifically, we sample 10k quad-chunks from the diffusion model, compute their surface point clouds (2048 points), and compare them to 10k sampled quad-chunks from the original scene.
4.3 Model and Baselines
We use triplane representations as a baseline for this task, given their widespread use in prior work [27, 47]. Note, we did not directly adopt BlockFusion’s triplane architecture, as it requires fitting triplanes for all dataset chunks, a process that took thousands of GPU hours on 3D-Front. Instead, we adopt an LRM [16] style VAE backbone (Sec.˜3.2).
We exclude dense grid methods [31, 33] due to their high memory demands, which scale cubically with resolution, compared to the quadratic growth of triplanes. In our task, the height can be up to about 8.5 times greater than the $x$ and $z$ dimensions for scene chunks in Figure˜5, further exacerbating memory usage. Additionally, LT3SD [33] requires multi-scale training for both the autoencoder and diffusion models, increasing training time.
VAE. For triplane baselines, we explore configurations with $V=3\times 4^{2}$, varying the normalization scale $S$, and the number of deconvolution layers, which affects output resolution (see Tab.˜2). For the vector set model, compressing to $V=16$ preserves geometric fidelity while enabling efficient diffusion training. Both representations use a channel size of $c=64$, and a sampled point cloud size of $N_{p}=4096$. We set the output triplane channel $c_{tri}=40$. The decoder has $L=24$ self-attention layers, and all models are trained for $160$ epochs. Each training iteration samples $4096$ query points per chunk for occupancy supervision.
Diffusion. For all models, we use a 25-layer UNet-like transformer [29] for diffusion. The triplane diffusion model operates on embeddings from the VAE with the output triplane resolution of $h_{tri}=w_{tri}=64$ and $S=6$. We train all models for $320$ epochs and a batch size of $192$. The transformer processes tokens of length $4\times V$ as the diffusion model generates a 2x2 grid of chunks at once.
4.4 Single Scene
In single-scene experiments, we demonstrate the effectiveness of representing scene chunks as vector sets compared to spatially structured latents like triplanes. VAE reconstruction results (Sec.˜4.4.1) and diffusion model comparisons (Sec.˜4.4.2) show that the vector set representation achieves better compression, leading to improved training efficiency and superior performance over triplanes.
4.4.1 VAE Reconstruction
Tab.˜3 presents the quantitative reconstruction results, showing that the vector set model outperforms all triplane settings. We also report whether occupancy queries are pruned using the predicted height $\hat{h}$ or the ground truth height $\tilde{h}$ when computing meshes for CD and F-Score. The similar values in both cases indicate the accuracy of our height prediction.
The triplane models are constrained by the VAE’s output resolution, with performance improving as resolution increases from $3\times 32^{2}$ to $3\times 64^{2}$. However, for tall scenes, coordinate clamping during triplane sampling prevents reconstruction of tall buildings exceeding the scaling factor of $S=6$. Increasing $S$ to $8.5$ (matching the tallest chunk) severely degrades performance due to precision loss from excessive compression along the y-axis. While further increasing resolution to $128^{2}$ via deconvolution is possible, it requires training on 4 L40S GPUs (Tab.˜2). In contrast, the vector set model provides better compression ($V=16$ vs $V=3\times 4^{2}$) while requiring similar training time and memory as the lowest-resolution triplane model, demonstrating its efficiency in representing scene chunks and superior performance. See additional results in the supplement.
@@ 4.4.1 VAE Reconstruction
Table 1: Comparison of VAE training resources for vector set vs triplane backbones. Training for most experiments was run on 2 L40S GPUs, total batch size and memory across all gpus are reported. The # Latents is the size of the latent for the VAE backbone and Output Res indicates the triplane size after deconvolution. Increasing the output triplane resolution to $3\times 128^{2}$ requires 4 GPUs for training.
@@
triplane 40 $\times 4^{2} 3×3$$\times 32^{2}$ 6 35.9 35.7
@@ 4.4.2 Diffusion
Figure 7: We compare generation results using RePaint [32] for outpainting with resampling steps $r=1$ and $r=5$.
@@
Figure 8: We present results from our NuiScene model trained on multiple scenes. In Figure (a), the top-right shows the generated geometry, while the top-left displays the textured model using SceneTex [6]. We render these models and feed them into Trellis [48] for reconstruction on the bottom (zoom in to see details). Our method produces better geometric detail. Figure (b) presents a large scene generated by our model, showcasing its ability to blend elements from different scenes in the dataset.
Quality. The FPD and KPD scores (Tab.˜4) show that the vector set diffusion model outperforms the baseline triplane diffusion model as well. Qualitative results for a 21x21 chunk scene (Figure˜6) demonstrate the vector set model’s ability to generate finer details, especially for buildings, whereas the triplane model struggles with details and introduces noisy artifacts which contributes to the higher FPD and KPD scores, likely due to compression of query points along the y axis. Additionally, as shown in Tab.˜2, the vector set model trains about 2.5 times faster and uses half the VRAM, benefiting from three times fewer tokens with a smaller $V$.
Table 5: We average times on a RTX 2080 Ti over 5 runs for generating $21\times 21$ chunks using RePaint with 5 resampling steps and our explicit outpainting method with the vecset diffusion model, and also report the triplane diffusion model’s generation time.
@@
vecset RePaint $21\times 21$ 1022.20 60.28
triplane explicit $1\times 21 22$1.68 30.84
vecset explicit $21\times 21$ 215.92 87.76
Outpainting. We compare against the inpainting method RePaint [32], widely used in prior works [47, 33, 27] for unbounded outpainting. Scenes are generated in the same order and with the same overlaps as our outpainting model. However, in RePaint, the mask and conditional embeddings are set to zero, with existing chunk conditions incorporated through its iterative re-sampling process. RePaint results for resampling steps $r=1$ and $r=5$ are shown in Figure˜7. Without resampling ($r=1$), inter-chunk connections lack coherence, and the process sometimes collapses into broken chunks, likely due to conditioning on noisy diffusion-generated chunks. With $r=5$, coherence improves, but the RePaint baseline produces less diverse scenes and occasional collapses (see supplement). Furthermore, re-sampling is significantly slower than our outpainting method (Tab.˜5), as inference time scales with the number of resampling steps, whereas our method does not require resampling. Note our outpainting model is parallel and can be potentially adapted into other methods [47, 33] for faster generation.
Our outpainting method is not perfect, and some chunks may still be poorly connected and contain noisy artifacts. In Figure˜6, the right orange box shows coherently connected chunks for our vector set diffusion model, while the left-side orange box highlights issues such as misaligned bridges or non-continuous chunks with visible gaps and seams.
@@ 4.5 Multiple Scenes
In Sec.˜4.4, we demonstrated the efficiency of vector sets, allowing handling of scene chunks with varying heights. As well as our outpainting model for fast large scale outdoor scene generation. However, training on a single scene leads to generated sub-scenes resembling the training data, despite the model’s ability to generate some novel compositions (see supplement). To assess its generalization ability and validate our curation process for NuiScene43, we extend training to three additional scenes using the vector set backbone.
Figure˜8 (b) shows a scene generated by our model, with purple boxes indicating how it integrates elements (rural houses, aqueduct, skyscraper) from two distinct scenes. The results demonstrates our model’s generalization ability, as such combinations are absent from training, and highlights the effectiveness of our curation process which unifies diverse scenes into a cohesive dataset, enabling the stitching of chunks from disjoint environments.
We further compare our approach with Trellis [48], a state-of-the-art image-to-3D object generation method, using TRELLIS-image-large without decimation. Figure˜8 (a) shows that compared to our model, Trellis produces coarser geometric details, as its single latent representation for entire scenes limits fine details in large environments. In contrast, our model represents each chunk separately, enabling higher-detail and unbounded scene generation. This highlights a key limitation of current object generation methods, which struggle to scale effectively for unbounded scene generation. See more results in the supplement.
5 Limitations
Currently, we are constrained to training on a limited number of scenes due to the high storage costs incurred by offline pre-sampling of chunks. More efficient online chunk sampling could enable training on our larger NuiScene43 dataset and improve inter-chunk consistency, potentially addressing issues where the model occasionally ignores conditions and generates discontinuous or misaligned chunks.
Another limitation is the lack of control or conditioning in our generative model, as we do not have labels for individual chunks. Leveraging foundation models [22, 1] for semi-automatic label generation of semantic maps or captions may be used for conditioning during training. A related challenge, shared by prior works [47, 33, 27, 31], is the absence of global context. The model only considers a limited local region, hindering large-scale planning. Exploring a global model capable of planning over large regions to mimic human-like design, such as coherent city layouts and road networks, is a promising direction for future work.
Finally, with the growing availability of 3D datasets [9], we aim to bridge the gap between methods like WonderWorld [51], which rely on 2D priors from T2I models for open-domain world generation. This mirrors the progression from DreamFusion [35], which distills T2I models for object-centric 3D generation, to models like TRELLIS that train directly on 3D data for faster and more consistent generation. While our work also leverages 3D supervision, the scale remains limited; as future work, we plan to expand our dataset further by semi-automating our curation pipeline. Nevertheless, we believe our dataset can serve as the basis of a benchmark for unbounded outdoor scene generation.
@@ 6 Conclusion
We introduce NuiScene, a method for generating expansive outdoor scenes across varied environments. Our model tackles key challenges such as large height variations and efficient compression of scene chunks using vector sets. In addition, our explicit outpainting model enables fast inference and improved coherence. To support this task, we curate NuiScene43, a dataset of moderate to large-scale scenes with varying styles, suitable for unified training.
@@
Figure 9: Additional generation results for our vector set diffusion model trained on 13 scenes. We show zoomed-in aerial shots on the top with street level views on the bottom respectively for the two scenes. The scenes shown here are of size $16\times 46$.
@@
Figure 10: Sample and depth mean deviation maps are calculated for sampling chunks from scenes.
@@
Figure 11: Three additional scenes used to train our 4-scene model. The top two sub-scenes are split from a large Objaverse scene for occupancy calculation. All scenes have fixed ground geometries, and their meshes are extracted via marching cubes on the occupancy grid.
Appendix A NuiScene43 Dataset
@@
We filter scenes from Objaverse to select 43 high quality moderately to large sized scenes (A.1). To ensure a unified scale between scenes, we annotate scenes with scales relative to each other, aligning them under a uniform scale (A.2). Finally, we preprocess the raw Objaverse mesh files for training by sampling point clouds, adjusting ground geometry, and converting them to occupancies (A.3). Furthermore, we describe the sampling maps used to sample quad chunks from scenes (A.4) as well as some dataset statistics (A.5). Please see our dataset page https://3dlg-hcvc.github.io/NuiScene43-Dataset for visualizations of all 43 scenes in NuiScene43.
A.1 Filtering
@@
We begin by filtering Objaverse [10] using multi-view embeddings from 12 object frames using DuoduoCLIP [25] . We use cosine similarity to query embeddings using text and images to retrieve scenes, then refine selection through manual labeling and neural network filtering trained on DuoduoCLIP embeddings, reducing 37k initial scenes to 2k. Since manually processing all scenes for scale labeling and ground alignment is impractical, we select 43 larger scenes for initial experiments.
A.2 Scale Labeling